OpenStudy (anonymous):

"the function f(x) is a difference of squares". how can I create a quadratic function with that?

OpenStudy (zzr0ck3r):

whats the question?

OpenStudy (anonymous):

What about it?

OpenStudy (anonymous):

I need to create a function using it.

OpenStudy (anonymous):

f(x)=20x+8

OpenStudy (anonymous):

how would I explain that to another person?

OpenStudy (anonymous):

okay lets use this example F(x)=x^2 , f(x) the x is the input , so lets pretend x is 2 , F(2) , that means to substitute 2 for every x in the equation so f(2)=2^2 simplified f(2)=4 , the input was 2 the output was 4

OpenStudy (ybarrap):

http://en.wikipedia.org/wiki/Difference_of_squares $$ a^2-b^2 $$ Now let a=x: $$ f(x)=x^2-b^2 $$ Which is a quadratic function that is a difference of squares

OpenStudy (anonymous):

Welcome to OpenStudy @reminisce ! Here's a link to help you around the site! http://openstudy.com/study#/updates/52f5a913e4b0512c88e5ef0f